OpenStudy (anonymous):

one card is selected from a standard deck of 52 cards. what is the probability that the card is either a club or jack?

OpenStudy (anonymous):

Let A=card drawn is a club B=card drawn is a jack . \[P(A \cup~B)=P \left( A \right)+P \left( B \right)-P \left( A~intersection~B \right)\] P(A)=13/52 P(B)=4/52 P( A intersection B)=1/52 find P(A U B)